OpenStudy (hitaro9):

Prove that if P>5 is a prime number, then P^2 +2 is never prime?

OpenStudy (hitaro9):

These prime proofs confuse me. Can someone help?

ganeshie8 (ganeshie8):

If you know fermat's little thm, for all \(p\ne 3\) we have: \[p^2 \equiv 1 \pmod{3}\] so \[p^2+2\equiv 1+2\equiv 3\equiv 0 \pmod{3}\] in other words \(3|(p^2+2)\) for all \(p \ne 3\)

OpenStudy (hitaro9):

I don't think we've seen that theorem. Is it difficult to prove without it?

ganeshie8 (ganeshie8):

Not at all, just show that \(p^2+2\) is divisible by 3 using division algorithm

ganeshie8 (ganeshie8):

By division algorithm any integer can be represented in one of the three forms : 3k, 3k+1, 3k+2 since p is a prime > 3, we only get two cases : p = 3k+1, 3k+2

OpenStudy (hitaro9):

Ohhh okay. That's pretty simple then.

ganeshie8 (ganeshie8):

case1 : p = 3k+1 p^2 + 2 = (3k+1)^2 + 2 = 9k^2+6k+1+2 = 3(3k^2+2k+1) which is divisible by 3

ganeshie8 (ganeshie8):

you can work the other case similarly :)

OpenStudy (hitaro9):

Yeah, that was a lot easier than I was thinking it was, Thanks a ton.

ganeshie8 (ganeshie8):

technically, you can prove ANY divisibility proofs using division algorithm. these get painful once the cases increase but you can see that these always work

OpenStudy (hitaro9):

Right, that makes sense. Again, thank you.